# Evidence-Focused Fact Summarization for Knowledge-Augmented Zero-Shot Question Answering

### v1

@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable zero-shot capabilities but often produce factual errors, known as hallucinations, particularly in knowledge-intensive tasks like Question Answering (QA). This happens because the static knowledge within LLM parameters may be incomplete, incorrect, or outdated, failing to keep pace with evolving real-world knowledge. Recent studies aim to remedy this by integrating external knowledge into LLMs Karpukhin et al. (2020b); Min et al. (2019). For instance, in open-domain QA frameworks, relevant passages are retrieved from document collections (e.g., Wiki corpus), which are then used to augment LLMs as contextual knowledge Ma et al. (2022a). By leveraging the external knowledge, LLMs are able to generate factual answers without the need to update their parameters.
Recently, there have been several attempts to consider Knowledge Graphs (KGs) as the knowledge source to augment LLMs for enhanced performance in knowledge graph QA (KGQA) Baek et al. (2023a); Wu et al. (2023); Sen et al. (2023). The primary challenge lies in the structured nature of KGs, which consist of facts presented in a triple format (head entity, relation, tail entity). Consequently, existing methods typically involve retrieving a set of relevant facts based on semantic similarity to a given question, followed by a process of verbalization into textual form for LLM prompting. Upon retrieving the relevant facts, one verbalization strategy is simply concatenating the facts in their triple-form text Baek et al. (2023a), while the other entails converting them into semantically-coherent textual description (i.e., free-form text) through the distillation of LLM’s ability to generate text from KGs Wu et al. (2023).
Despite their remarkable efficacy, existing verbalization strategies for providing contextual knowledge exhibit critical limitations. (1) Low density of evidence: Both concatenation and linearization of the facts Baek et al. (2023a); Wu et al. (2023) are highly likely to include duplicated entities or relations due to limited flexibility; this eventually degrades the density of useful evidence within the contextual knowledge, hindering the ability to answer questions effectively. (2) Low clarity of evidence: While contextual knowledge describes factual information, they often fail to highlight the evidence necessary for answering questions. This lack of focus can lead to noise from irrelevant facts, which can detrimentally impact the LLM’s ability to provide accurate answers. For this reason, even when sufficient evidence (or an exact answer span) is presented in the knowledge context, LLMs may struggle to properly identify it.
@@ 2.1 KG-Augmented LLM Prompting for QA
We focus on a QA approach that leverages LLMs’ zero-shot capability for answering the question, enhanced with external knowledge from KGs.111 In this work, we focus on symbolic facts as external knowledge, and the terms “facts” and “triples” are used interchangeably in that a single fact has the form of triple, i.e., (head entity, relation, tail entity). This approach mainly consists of the three stages, referred to as retrieve-verbalize-inject Baek et al. (2023a); Wu et al. (2023), each of which is responsible for (1) fact retrieval, (2) fact verbalization, and (3) fact injection, respectively.
Fact retrieval from knowledge graph. The first stage aims to retrieve question-associated facts from KGs via entity linking, and then to select only top- ones based on their semantic relevance to the input question. The process of knowledge retrieval begins with linking the question to corresponding entities within the KG, which involves span detection, entity disambiguation, and relation classification. Following this, pertinent facts are collected, by retrieving all neighbor triples of the linked entities Baek et al. (2023a), or sampling triples from predicted relation paths Wu et al. (2023). To select only the most relevant facts directly beneficial for the QA task, recent studies utilize semantic similarities between each fact and the question, employing either a pretrained sentence encoder Song et al. (2020); Karpukhin et al. (2020a); Xiong et al. (2020) or one fine-tuned specifically for direct fact retrieval Baek et al. (2023b).
Fact verbalization into various form text. The second stage is fact verbalization, which refers to the task of transforming symbolic facts into textual strings, for feeding them into the LLM as the contextual knowledge. The linear verbalization simply concatenates the head, relation, and tail texts in the triple while keeping the structured format (i.e., triple-form text) Baek et al. (2023a), or use manually-designed templates and heuristics for linearization Oguz et al. (2022); Ma et al. (2022b). On the other hand, the graph-to-text verbalization transforms the input facts into the plausible and coherent text by using a fine-tuned model Ribeiro et al. (2021) or prompting LLMs Wu et al. (2023).
Fact injection for question answering. The last stage is prompting the LLM to generate the answer, while incorporating the verbalized facts as contextual knowledge. This process, also known as knowledge-augmented LLM prompting for zero-shot QA, gathers the output as the predicted answer. To address various scenarios where the contextual knowledge provides sufficient or insufficient evidence to answer the question, detailed instructions are explicitly provided to allow the LLM to utilize its internal knowledge if it determines that the contextual knowledge is insufficient for inferring the answer. The prompts are in Appendix C.2.
@@ 2.2 Analysis on Verbalized Facts
In Figure 2 Lower, linear verbalization tends to scatter obvious evidence (i.e., answer span) randomly within the contextual knowledge. Their placement seems to rely on the rank obtained from fact retrieval, rather than being positioned at the forefront of the verbalized texts for emphasis. Moreover, the average semantic similarity 222Average semantic similarity is calculated through the cosine similarity between embeddings encoded via MPNet. of their verbalized facts with the question is lower compared to other methods. This shows that their outputs may include noisy or irrelevant information, or they might not clearly highlight semantically relevant evidence. In essence, clarity of evidence is not adequately addressed in this approach.
@@ 3.2 LLM Fine-Tuning for EFSum
To enhance the quality of summaries, we propose EFSum, a fact verbalization model based on an open-source LLM,333In this work, we choose Llama2-7B Touvron et al. (2023) as the backbone open-source LLM for EFSum. specifically fine-tuned for evidence-focused fact summarization. For helpful and faithful summary generation, EFSum is optimized in two steps: (1) LLM distillation and (2) preference alignment. Figure 3 illustrates the overall EFSum framework.
@@ 3.2.2 Alignment with Summary Preference
where . By optimizing the model using preferred-dispreferred summary pairs, our obtained model is trained to be more biased towards helpful and faithful summary while avoiding unhelpful or unfaithful summary . Note that has been specifically trained for each QA model, as there are different preferences regarding summary helpfulness across various QA models.
@@ 4 Experiments
=200 |
=400 |
=200 |
=400 |
=200 |
=400 |
| WebQSP | No knowledge |
0.506 |
0.506 |
0.409 |
0.409 |
0.539 |
0.539 |
| KAPING Baek et al. (2023a) |
0.507 |
0.538 |
0.391 |
0.439 |
0.517 | 0.519 | |
| KG2Text Ribeiro et al. (2021) |
0.476 |
0.476 |
0.316 |
0.321 |
0.439 |
0.481 |
| Rewrite Wu et al. (2023) |
0.444 |
0.525 |
0.350 |
0.431 |
0.462 |
0.511 | |
| EFSum (Ours) | 0.537 | 0.538 | 0.447 | 0.468 |
0.457 |
0.491 |
| EFSum (Ours) | 0.559 | 0.569 | 0.458 | 0.500 | 0.489 |
0.497 |
| Mintaka | No knowledge |
0.557 |
0.557 |
0.211 |
0.211 |
0.492 |
0.492 |
| KAPING Baek et al. (2023a) | 0.532 | 0.550 |
0.269 |
0.299 |
0.463 | 0.481 | |
| KG2Text Ribeiro et al. (2021) |
0.425 |
0.406 |
0.220 |
0.224 |
0.462 |
0.448 |
| Rewrite Wu et al. (2023) |
0.525 |
0.535 |
0.287 |
0.322 |
0.415 |
0.459 |
| EFSum (Ours) |
0.486 |
0.433 |
0.318 |
0.321 |
0.459 |
0.455 |
| EFSum (Ours) | 0.538 |
0.504 |
0.327 | 0.327 | 0.474 | 0.486 | |
@@
Random |
Popular |
MPNet |
Random |
Popular |
MPNet |
Random |
Popular |
MPNet |
| WebQSP | No knowledge |
0.506 |
0.506 |
0.506 |
0.409 |
0.409 |
0.409 |
0.539 |
0.539 |
0.539 |
| KAPING Baek et al. (2023a) |
0.441 |
0.437 |
0.538 |
0.297 |
0.329 |
0.439 |
0.476 | 0.490 | 0.519 | |
| KG2Text Ribeiro et al. (2021) |
0.469 |
0.468 |
0.476 |
0.317 |
0.276 |
0.321 |
0.465 |
0.451 |
0.481 |
| Rewrite Wu et al. (2023) |
0.473 |
0.445 |
0.525 |
0.323 |
0.348 |
0.431 |
0.458 |
0.439 |
0.511 | |
| EFSum (Ours) | 0.542 | 0.534 | 0.538 | 0.443 | 0.442 | 0.468 | 0.477 |
0.472 |
0.491 |
| EFSum (Ours) | 0.475 | 0.539 | 0.569 | 0.500 | 0.505 | 0.500 |
0.457 |
0.488 |
0.497 |
| Mintaka | No knowledge |
0.557 |
0.557 |
0.557 |
0.211 |
0.211 |
0.211 |
0.492 |
0.492 |
0.492 |
| KAPING Baek et al. (2023a) | 0.541 |
0.479 |
0.550 |
0.215 |
0.191 |
0.299 |
0.467 |
0.463 |
0.492 | |
| KG2Text Ribeiro et al. (2021) |
0.411 |
0.43 |
0.406 |
0.224 |
0.219 |
0.224 |
0.457 |
0.459 |
0.481 |
| Rewrite Wu et al. (2023) | 0.522 | 0.487 | 0.535 |
0.257 |
0.264 |
0.322 |
0.414 |
0.447 |
0.448 |
| EFSum (Ours) |
0.469 |
0.467 |
0.433 |
0.293 | 0.297 |
0.321 |
0.464 | 0.472 |
0.459 |
| EFSum (Ours) |
0.470 |
0.502 |
0.504 |
0.301 | 0.323 | 0.327 |
0.448 |
0.469 | 0.501 | |
@@ 4.1 Experimental Settings
Datasets. WebQuestionsSP (WebQSP) Yih et al. (2016) is a KGQA dataset that filters questions from the WebQuestions Berant et al. (2013) dataset to include only those answerable via Freebase, and provides SPARQL queries for them. For convenience, we use WebQSP-WD Sorokin and Gurevych (2018), in which each question from WebQSP is pre-linked to the Wikidata KG. We use a test set comprising 1,033 examples for evaluation. Mintaka Sen et al. (2022) is a QA dataset that encompasses eight different complexity types. To mainly focus on the efficacy of fact verbalization methods for zero-shot QA, we use only 795 “generic” type questions out of 4,000 in the test set for evaluation, similar to previous work Razzhigaev et al. (2023).
@@
Relevant fact retrieval. In retrieving question-related facts, we exclusively utilize triples composed of the one-hop neighbors of the question entity. However, these one-hop triples contain information that may not always be relevant. To address this issue, we employ MPNet Song et al. (2020) to retrieve only the top- triples with the highest semantic similarity to the question representation, following Baek et al. (2023a). When calculating semantic similarity, we use the linear verbalization approach, which involves combining the subject, relation, and object texts from the triple.
@@
=10 |
=30 |
=10 |
=30 |
=10 |
=30 |
| WebQSP | No knowledge |
0.617 |
0.607 |
0.498 |
0.451 |
0.646 |
0.628 |
| KAPING Baek et al. (2023a) |
0.777 |
0.771 | 0.643 | 0.738 | 0.668 | 0.699 | |
| KG2Text Ribeiro et al. (2021) |
0.589 |
0.608 |
0.467 |
0.457 |
0.409 |
0.536 |
| Rewrite Wu et al. (2023) |
0.628 |
0.728 |
0.533 |
0.664 |
0.594 |
0.688 | |
| EFSum (Ours) | 0.788 |
0.755 |
0.629 |
0.711 |
0.497 |
0.571 |
| EFSum (Ours) | 0.786 | 0.783 | 0.644 | 0.741 | 0.599 |
0.666 |
| Mintaka | No knowledge |
0.737 |
0.755 |
0.340 |
0.308 |
0.622 |
0.625 |
| KAPING Baek et al. (2023a) |
0.816 |
0.904 | 0.645 | 0.746 | 0.702 | 0.772 | |
| KG2Text Ribeiro et al. (2021) | 0.840 |
0.663 |
0.724 |
0.500 |
0.756 |
0.712 |
| Rewrite Wu et al. (2023) | 0.842 | 0.908 |
0.640 |
0.759 |
0.684 |
0.750 |
| EFSum (Ours) |
0.732 |
0.829 |
0.662 |
0.737 |
0.781 | 0.807 | |
| EFSum (Ours) |
0.768 |
0.851 |
0.605 |
0.732 |
0.702 |
0.768 |
@@ 4.2 Effectiveness of Dense Evidence (RQ1)
Effect of knowledge augmentation. First of all, we observe that knowledge augmentation for zero-shot QA does not always produce positive results. Knowledge augmentation cannot be helpful in two scenarios: where the model’s ability to ground input knowledge is lacking, or where the retrieved knowledge is noisy while the QA model’s internal knowledge is sufficient. In Table 1, when Llama2-7B-Chat was used as the QA model, it demonstrates higher performance under the “No knowledge” condition across both datasets (i.e., WebQSP and Mintaka) compared to other baselines. This is indicative of Llama2-7B-Chat’s limited capability in utilizing the provided knowledge. Moreover, in case of GPT-3.5-turbo on Mintaka, the “No knowledge” condition exhibits superior performance over other baselines. This is because the Mintaka dataset comprises more specific and challenging questions compared to WebQSP (e.g., “What competition did Dean Koontz win as a senior in college?”). For such questions, the correct answer often does not feature within the retrieved facts, suggesting that introducing noisy knowledge can be counterproductive. Furthermore, it is also due to GPT-3.5-turbo having sufficient internal knowledge.
@@
Compatibility with various retrievers. We investigate the QA performance using different types of fact retrieval, to assess the robustness across various knowledge qualities: randomly selected knowledge (Random), the knowledge possessing the most frequently occurring relation (Popular), and question-relevant knowledge (MPNet). In Table 2, EFSum achieves the highest accuracy across most datasets and QA models. These results imply the superiority of EFSum, regardless of the retriever used. Considering the robustness demonstrated with basic retrievers, we expect that the performance of EFSum can be further enhanced with more sophisticated retrievers.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable zero-shot abilities but often produce factual errors, known as hallucinations, particularly in knowledge-intensive tasks like Question Answering (QA). This happens because the static knowledge within LLM parameters may be incomplete, incorrect, or outdated, failing to keep pace with evolving real-world knowledge. Recent studies remedy this by integrating external knowledge into LLMs Karpukhin et al. (2020b); Min et al. (2019).
As one form of external knowledge, Knowledge Graphs (KGs) have been considered as the knowledge source to augment LLMs for enhanced performance in knowledge graph QA (KGQA) Baek et al. (2023a); Wu et al. (2023); Sen et al. (2023). The key challenge of utilizing KGs, which consist of a set of (head entity, relation, tail entity) triples, is to bridge the modality gap between graphs and text. Efforts on connecting the gap mostly fall into either training additional layers to blend two representations, or verbalizing graphs into texts. While training fusion layers Yasunaga et al. (2022a, b) can make models expressive on two different modalities, it takes expensive computaions, and needs to be trained when KGs are updated. On the other hand, recent studies proposed verbalizing KGs into text-form, without restirctions of trainining LLMs for QA. For instance, one strategy is simply concatenating the facts in their triple-form text Baek et al. (2023a), and another entails converting them into semantically-coherent textual description (i.e., free-form text) through the distillation of LLM’s ability to generate text from KGs Wu et al. (2023).
Despite their remarkable efficacy, existing verbalization strategies for providing contextual knowledge exhibit critical limitations. (1) Low density of evidence: Both concatenation and linearization of the facts Baek et al. (2023a); Wu et al. (2023) are highly likely to include duplicated entities or relations due to limited flexibility; this eventually degrades the density of useful evidence within the contextual knowledge, hindering the ability to answer questions effectively. (2) Low clarity of evidence: While contextual knowledge describes factual information, they often fail to highlight the evidence necessary for answering questions. This lack of focus can lead to noise from irrelevant facts, which can detrimentally impact the LLM’s ability to provide accurate answers.
@@
For reproducibility, our codes are publicly available at https://github.com/anon809/EFSum
@@ 2.1 KG-Augmented LLM Prompting for QA
We focus on a QA approach that leverages LLMs’ zero-shot capability for answering the question, enhanced with external knowledge from KGs.
Fact retrieval from knowledge graph. The first stage aims to retrieve question-associated facts from KGs via entity linking, and then to select only top- ones based on their semantic relevance to the input question. To select only the most relevant facts, recent studies utilize semantic similarities between each fact and the question, employing either a pretrained sentence encoder Karpukhin et al. (2020a); Xiong et al. (2020) or one fine-tuned specifically for direct fact retrieval Baek et al. (2023b). In this work, we utilized the former strategy Song et al. (2020), if not stated.
Fact verbalization into various form text. The second stage is fact verbalization, which refers to the task of transforming symbolic facts into textual strings, for feeding them into the LLM as the contextual knowledge. The linear verbalization simply concatenates the head, relation, and tail texts in the triple while keeping the structured format (i.e., triple-form text) Baek et al. (2023a), or use manually-designed templates and heuristics for linearization Oguz et al. (2022); Ma et al. (2022). On the other hand, the graph-to-text verbalization transforms the input facts into the plausible and coherent text by using a fine-tuned model Ribeiro et al. (2021) or prompting LLMs Wu et al. (2023).
Fact injection for question answering. The last stage is prompting the LLM to generate the answer with the verbalized facts. This process, also known as knowledge-augmented LLM prompting for zero-shot QA, gathers the output as the predicted answer. To handle with insufficient evidence of knowledge, the detailed instructions are provided to allow the LLM to utilize its internal knowledge if needed. The prompts are in Appendix C.2.
@@ 2.2 Analysis on Verbalized Facts
In Figure 2 Lower, the linear verbalization tends to scatter obvious evidence (i.e., answer span) randomly within the contextual knowledge. Their placement seems to rely on the rank obtained from fact retrieval, rather than being positioned at the forefront of the verbalized texts for emphasis. Moreover, the average semantic similarity 111Average semantic similarity is calculated through the cosine similarity between embeddings encoded via MPNet. of their verbalized facts with the question is lower compared to other methods. This shows that their outputs may include noisy or irrelevant information, or they might not clearly highlight semantically relevant evidence. In essence, clarity of evidence is not adequately addressed in this approach.
@@ 3.2 LLM Fine-Tuning for EFSum
To enhance the quality of summaries, we propose EFSum, a fact verbalization model based on an open-source LLM,222In this work, we choose Llama2-7B Touvron et al. (2023) as the backbone open-source LLM for EFSum. specifically fine-tuned for evidence-focused fact summarization. For helpful and faithful summary generation, EFSum is optimized in two steps: (1) LLM distillation and (2) preference alignment. Figure 3 illustrates the overall EFSum framework.
@@ 3.2.2 Alignment with Summary Preference
where . By optimizing the model using preferred-dispreferred summary pairs, our obtained model is trained to be more biased towards helpful and faithful summary while avoiding unhelpful or unfaithful summary . Note that has been specifically trained for each QA model, as there are different preferences regarding summary helpfulness across various QA models. The following is an example of an output fact summary.
Question: where was george washington carver from?
Answer: Diamond
Facts: (George Washington Carver, occupation, biologist),(George Washington Carver, interested in, botany),(George Washington Carver, occupation, university teacher),(George Washington Carver, place of birth, Diamond),(George Washington Carver, given name, George), (George Washington Carver, field of work, pedagogy),(George Washington Carver, relative, Moses Carver),(George Washington Carver, place of birth, United States of America),(George Washington Carver, residence, Tuskegee),(George Washington Carver, occupation, inventor)
Summary: George Washington Carver, the renowned biologist, was born in Diamond, United States of America. He took great interest in botany and became a university teacher, focusing on pedagogy. Additionally, Carver contributed significantly as an inventor. His relative, Moses Carver, also had a close association with him. Carver later resided in Tuskegee.
@@ 4 Experiments
=200 | =400 | =200 | =400 | =200 | =400 | ||
WebQSP | No knowledge | 0.506 | 0.506 | 0.409 | 0.409 | 0.539 | 0.539 |
KAPING Baek et al. (2023a) | 0.507 | 0.538 | 0.391 | 0.439 | 0.517 | 0.519 | |
KG2Text Ribeiro et al. (2021) | 0.476 | 0.476 | 0.316 | 0.321 | 0.439 | 0.481 | |
Rewrite Wu et al. (2023) | 0.444 | 0.525 | 0.350 | 0.431 | 0.462 | 0.511 | |
EFSum (Ours) | 0.537 | 0.538 | 0.447 | 0.468 | 0.457 | 0.491 | |
EFSum (Ours) | 0.559 | 0.569 | 0.458 | 0.500 | 0.489 | 0.497 | |
Mintaka | No knowledge | 0.540 | 0.540 | 0.228 | 0.228 | 0.440 | 0.440 |
KAPING Baek et al. (2023a) | 0.539 | 0.539 | 0.269 | 0.279 | 0.402 | 0.407 | |
KG2Text Ribeiro et al. (2021) | 0.492 | 0.491 | 0.234 | 0.234 | 0.377 | 0.378 | |
Rewrite Wu et al. (2023) | 0.515 | 0.521 | 0.280 | 0.288 | 0.394 | 0.386 | |
EFSum (Ours) | 0.496 | 0.491 | 0.312 | 0.321 | 0.423 | 0.418 | |
EFSum (Ours) | 0.474 | 0.449 | 0.326 | 0.338 | 0.405 | 0.406 |
@@
Random | Popular | MPNet | Random | Popular | MPNet | Random | Popular | MPNet | ||
WebQSP | No knowledge | 0.506 | 0.506 | 0.506 | 0.409 | 0.409 | 0.409 | 0.539 | 0.539 | 0.539 |
KAPING Baek et al. (2023a) | 0.441 | 0.437 | 0.538 | 0.297 | 0.329 | 0.439 | 0.476 | 0.490 | 0.519 | |
KG2Text Ribeiro et al. (2021) | 0.469 | 0.468 | 0.476 | 0.317 | 0.276 | 0.321 | 0.465 | 0.451 | 0.481 | |
Rewrite Wu et al. (2023) | 0.473 | 0.445 | 0.525 | 0.323 | 0.348 | 0.431 | 0.458 | 0.439 | 0.511 | |
EFSum (Ours) | 0.542 | 0.534 | 0.538 | 0.443 | 0.442 | 0.468 | 0.477 | 0.472 | 0.491 | |
EFSum (Ours) | 0.475 | 0.539 | 0.569 | 0.500 | 0.505 | 0.500 | 0.457 | 0.488 | 0.497 | |
Mintaka | No knowledge | 0.540 | 0.540 | 0.540 | 0.228 | 0.228 | 0.228 | 0.440 | 0.440 | 0.440 |
KAPING Baek et al. (2023a) | 0.553 | 0.516 | 0.539 | 0.201 | 0.198 | 0.279 | 0.417 | 0.398 | 0.407 | |
KG2Text Ribeiro et al. (2021) | 0.505 | 0.500 | 0.492 | 0.220 | 0.235 | 0.234 | 0.421 | 0.389 | 0.378 | |
Rewrite Wu et al. (2023) | 0.527 | 0.524 | 0.515 | 0.230 | 0.224 | 0.288 | 0.393 | 0.374 | 0.386 | |
EFSum (Ours) | 0.454 | 0.492 | 0.496 | 0.213 | 0.215 | 0.321 | 0.390 | 0.392 | 0.418 | |
EFSum (Ours) | 0.427 | 0.425 | 0.474 | 0.292 | 0.243 | 0.338 | 0.397 | 0.393 | 0.406 |
@@ 4.1 Experimental Settings
Datasets. WebQuestionsSP (WebQSP) Yih et al. (2016) is a KGQA dataset that filters questions from the WebQuestions Berant et al. (2013) dataset to include only those answerable via Freebase, and provides SPARQL queries for them. For convenience, we use WebQSP-WD Sorokin and Gurevych (2018), in which each question from WebQSP is pre-linked to the Wikidata KG. We use a test set comprising 1,033 examples for evaluation. Mintaka Sen et al. (2022) is a QA dataset that encompasses eight different complexity types. Most question-answer pairs can only be solved by utilizing multi-hop reasoning or the attributes of multiple entities. We use a test set of Mintaka which has 4,000 examples for evaluation.
@@
Relevant fact retrieval. In retrieving question-related facts, as described in the Section 2.1, we employ MPNet Song et al. (2020) to retrieve only the top- triples among given KG, with the highest semantic similarity to the question representation, following Baek et al. (2023a). Since processing entire KGs is impractical, we focus on retrieving information from the -hop neighbors of question entities within a KG. The value of is given by the specific KGQA dataset to answer the question. In our experiment, is set to 1 for WebQSP and 2 for Mintaka. When calculating semantic similarity, we use the linear verbalization approach, which involves combining the subject, relation, and object texts from the triple.
@@
=10 | =30 | =10 | =30 | =10 | =30 | ||
WebQSP | No knowledge | 0.617 | 0.607 | 0.498 | 0.451 | 0.646 | 0.628 |
KAPING Baek et al. (2023a) | 0.777 | 0.771 | 0.643 | 0.738 | 0.668 | 0.699 | |
KG2Text Ribeiro et al. (2021) | 0.589 | 0.608 | 0.467 | 0.457 | 0.409 | 0.536 | |
Rewrite Wu et al. (2023) | 0.628 | 0.728 | 0.533 | 0.664 | 0.594 | 0.688 | |
EFSum (Ours) | 0.788 | 0.755 | 0.629 | 0.711 | 0.497 | 0.571 | |
EFSum (Ours) | 0.786 | 0.783 | 0.644 | 0.741 | 0.599 | 0.666 | |
Mintaka | No knowledge | 0.810 | 0.788 | 0.492 | 0.444 | 0.783 | 0.719 |
KAPING Baek et al. (2023a) | 0.912 | 0.869 | 0.723 | 0.673 | 0.832 | 0.808 | |
KG2Text Ribeiro et al. (2021) | 0.879 | 0.768 | 0.536 | 0.491 | 0.799 | 0.727 | |
Rewrite Wu et al. (2023) | 0.901 | 0.875 | 0.720 | 0.691 | 0.843 | 0.792 | |
EFSum (Ours) | 0.920 | 0.887 | 0.742 | 0.735 | 0.849 | 0.806 | |
EFSum (Ours) | 0.893 | 0.824 | 0.745 | 0.719 | 0.852 | 0.826 |
@@ 4.2 Effectiveness of Dense Evidence (RQ1)
Effect of knowledge augmentation. First of all, we observe that knowledge augmentation for zero-shot QA does not always produce positive results. Knowledge augmentation cannot be helpful in two scenarios: where the model’s ability to ground input knowledge is lacking, or where the retrieved knowledge is noisy while the QA model’s internal knowledge is sufficient. In Table 1, when Llama2-7B-Chat was used as the QA model, it demonstrates higher performance under the “No knowledge” condition across both datasets (i.e., WebQSP and Mintaka) compared to other baselines. This is indicative of Llama2-7B-Chat’s limited capability in utilizing the provided knowledge.
@@
Compatibility with various retrievers. We investigate the QA performance using different types of fact retrieval, to assess the robustness across various knowledge qualities: randomly selected knowledge (Random), the knowledge possessing the most frequently occurring relation (Popular), and question-relevant knowledge (MPNet). In Table 2, EFSum achieves the highest accuracy across most datasets and QA models, regardless of the retriever used, which implies that our method can exact useful facts from noisy input triples.